Box office sales for movie releases are widely reported and discussed. Find out (approximately) what percentage of box office sales in the United States go back to the movie's distributor. Then calculate how much profit a movie that cost $250 million to produce and $200 million to market and distribute generates if it has a domestic box office of $700 million. Assume no foreign box office or any other sources of cash or any other expenses, and that there are no time-value of money effects to account for. Enter your answer in millions of dollars (e.g., enter 20 for $20 million, or -50 for a $50 million loss).
Answer #1

Answer- Total cost to produce a movie = 250+200

=450

proportionate share of cost by movie distributor = 200/450

=44.44%

Revenue earned by movie distributor

Total sales (revenue) generated = 700

The distribution of revenue is done on proportionate basis

= 700*44.44%

= 311.11

Profit earned

Profit = Sales - Cost

=700-450

=250

Profit Percentage

= (profit / cost)*100

=(250/450)*100

=55.55%

Profit earned by movie distributor

= 250*44.44%

=111.11